OpenStudy (anonymous):

Passes through (3,1) and perpendicular to graph of y=1/2x+3? Explain. Passes through (3,1) and perpendicular to graph of y=1/2x+3? Explain. @Mathematics

OpenStudy (anonymous):

Perpendicular slopes will equal -1 so to find the perpendicular slope of 1/2 you do this \[\frac{1}{2} + x = -1\] Solve for x and you get the perpendicular slope of -2

OpenStudy (anonymous):

now that you have a point on the graph and a slope, use the point slope formula

OpenStudy (anonymous):

\[y - 1 = -2(x - 3)\] solve for y \[y = -2x + 7\]

OpenStudy (anonymous):

Thanks, I just don't understand this stuff.. My teacher is WACK and won't explain it to me.. Thank you though!

OpenStudy (anonymous):

Just remember that perpendicular slopes add up to -1.. or the easy way is to do the negative reciprocal. Reciprocal means to flip the fraction and then all you have to do is negate it. So if you have a slope of 3/2 and you want the perpendicular slope all you do is flip the fraction... 2/3.. then negate it.. so the slope perpendicular to 3/2 would be -2/3. Understand??

OpenStudy (anonymous):

Yeah I think I got it! My teacher just needs to teach it like that! THANKS!